# Thermal one-loop self-energy correction for hydrogen-like systems: relativistic approach

### v1

@@ III Technical details, numerical results and discussion
Specifically, the values given in the Table 2 allow for an unambiguous determination of the influence of the Lamb shift on the thermal shifts of hydrogen atomic levels. It can be seen that the contribution does not exceed a tenth of a percent of the total shift. In turn, the values show a fundamental difference between two approaches: a fully relativistic one that takes into account the Lamb shift (this work), and an approximate non-relativistic [farley].
@@
Being the most accurately calculated to date for the hydrogen atom, the values for the thermal shift, presented in Table 2, can be used to determine the thermal frequency shifts in transitions between fine sublevels. In accordance with the quantum mechanical formalism for incorporating thermal radiation corrections to atomic levels [farley], or via the QED one-loop perturbative framework [PRA_2015, Lopez_2025], the dominant thermal contribution arises due to the static dipole polarizability. Considering that in this case the perturbation operator does not act on the spin part of the electronic function, such contributions cancel out for states with equal principal quantum number $n$ and orbital angular momentum $l$, but different total angular momenta $j$.
@@
The detailed theoretical analysis assigning the influence of the "electric component"  (Stark effect) and the "magnetic component"  (Zeeman effect), as well as the derivation of other thermal corrections, has been presented quite recently in the work [Lopez_2025]. A distinctive feature of the approach presented in this study is that the fine structure of atomic levels is automatically taken into account within the framework of Dirac’s theory. Thus, by performing calculations with a sufficient number of significant digits, we can accurately account for the mentioned reduction.
@@
Below we present a series of values for the hydrogen atom and several hydrogen-like ions for the 1s2s1s-2s and 1s3s1s-3s transition frequencies as a function of temperature. It is well known that the Stark thermal shift for certain transition frequencies can vanish at the magic temperature [PhysRevA.100.023417]. On the other hand, such a cancellation of the leading order shift should be verified for the possibility of residual contributions (in our case, relativistic effects [Lopez_2025]). The results of calculations for temperatures corresponding to laboratory conditions are summarized in Table 4, which also provides the values of $T_{m}$ at which the thermal frequency shift vanishes.
@@ H $\Re\Delta\varepsilon_{3s_{1/2}}^{\beta}-\Re\Delta\varepsilon_{1s_{1/2}}^{\beta}$ $T_{m}$
To perform the numerical calculations, see Eq. (13), the values for the hyperfine splitting were borrowed from Refs. [HFS_Shabaev_1994, HH-tab, HFS_He+, PhysRevA.108.052802, PhysRevA.52.3686]. First, what follows from Table 5 is the cubic decrease of the thermal frequency shift with an increase in the principal quantum number (easily verifiable for $s$-states). Secondly, it can be noted that the thermal shift actually increases cubically with the growth of the nuclear charge $Z$, as well as quadratically with the increase in temperature approaching laboratory conditions.
The corresponding estimate arises from formula (13), considering that at room temperature, the thermal radiation frequency $\omega$ is on the order of $2.82\beta\approx 0.00268$ in atomic units, while for hydrogen $\Delta\varepsilon_{\rm HFS}\approx 2.158\times 10^{-7}$ a.u. Then, expanding expression (13) in a series with respect to the small parameter $\Delta\varepsilon_{\rm HFS}$, one can obtain $\Delta\varepsilon_{\rm HFS}^{\beta}\sim\Delta\varepsilon_{\rm HFS}(k_{\rm B}T)^{2}$, $\Delta\varepsilon_{\rm HFS}\sim Z^{3}$ [HFS_Shabaev_1994, Grespo-hfs]. However, this estimate is valid only up to $Z=23$. For larger values of the nuclear charge, it is necessary to perform a decomposition by the frequency of photons $\omega$ from thermal radiation. Then, a different dependence will emerge, and moreover, one of the opposite sign: $\Delta\varepsilon_{\rm HFS}^{\beta}\sim-(k_{\rm B}T)^{4}/\Delta\varepsilon_{\rm HFS}$. In the intermediate region, the complete form of equation (13) should be considered. In certain specific cases, the effect of finite lifetime can play a role [Lopez_2025].
@@
The value of the thermal shift to the hyperfine transition frequency in the hydrogen atom agrees well with the results of [Itano], but it is negligibly small even with respect to the accuracy of the experiment of [Hellwig]. However, for neutral and singly ionized atoms, the thermal shift of the hyperfine splitting remains the same in terms of the ratio of the shift to the transition frequency and undoubtedly makes a significant contribution to the budget of experimental error [PhysRevA.110.043108, Lopez_2025]. Below we present a graph in Fig. 6 illustrating the $Z$-dependence of the shift $\Delta\varepsilon_{a}^{\beta}$ expressed by Eq. (13).
Figure 6: Thermal frequency shift of the hyperfine transitions for the ground state of the hydrogen atom and hydrogen-like ions as a function of nuclear charge $Z$. The dashed lines represent the asymptotic estimate $Z^{3}$ and $1/Z^{3}$, the values of frequency shift are given in eV.
@@
Figure 7: The graph shows the dependence of the hyperfine splitting interval of the ground state in hydrogen-like ions on the nuclear charge $Z$ (in logarithmic scale). The points represent the used values, $\Delta\varepsilon_{\rm HFS}$ in eV. The dashed line corresponds to the dependence $\frac{1}{2}\Delta\varepsilon_{\rm HFS}^{Z=1}Z^{3}$, chosen as the interpolation according to [HFS_Shabaev_1994], see Table 7 for details.
@@
Table 7: Thermal shift for hyperfine transition frequencies of the ground state as a function of nuclear charge $Z$. The energy intervals used are indicated in the second column, and the frequency thermal shift is in the third column. All values are given in eV at a temperature of $T=300$ K.
@@
Combining the results from Tables 5, 6, and Fig. 6, one can conclude that the thermal shift for the hyperfine splitting frequency of the ground state (despite the cubic increase with $Z$) is unlikely to be related to corresponding measurements in highly charged ions, see, e.g., [Grespo-hfs].

### v2

@@ III Technical details, numerical results and discussion
Specifically, the values given in the Table 2 allow for an unambiguous determination of the influence of the Lamb shift on the thermal shifts of hydrogen atomic levels. It can be seen that the contribution does not exceed a tenth of a percent of the total shift. In turn, the values show a fundamental difference between two approaches: a fully relativistic one that takes into account the Lamb shift (this work), and an approximate non-relativistic [14].
[…]
black In addition to Table 2, we have performed calculations of the thermal one-loop correction for several hydrogen-like ions at various temperatures for low-lying states. The results are collected in Table 3.
[…]
$Z\quad$ State 33 K 55 K $10$ K 2020 K $30$ K $50$ K $100$ K $200$ K
11 $1s_{1/2}$ $4.3(4)\times 10^{-10}$ $3.1(1)\times 10^{-9}$ $4.83(5)\times 10^{-8}$ $7.67(2)\times 10^{-7}$ $3.879(4)\times 10^{-6}$ $2.991(1)\times 10^{-5}$ $4.7842(5)\times 10^{-4}$ $7.6545(2)\times 10^{-3}$
$2s_{1/2}$ $4.8(1)\times 10^{-6}$ $1.24(1)\times 10^{-5}$ $5.08(2)\times 10^{-5}$ $1.897(5)\times 10^{-4}$ $3.724(8)\times 10^{-4}$ $5.32(1)\times 10^{-4}$ $7.416(5)\times 10^{-3}$ $1.82900(3)\times 10^{-1}$
$\it 4.58(6)\times 10^{-6}$} $\it 1.20(3)\times 10^{-5}$} $\it 4.93(5)\times 10^{-5}$} $\it 1.747(8)\times 10^{-4}$} $\it 3.42(1)\times 10^{-4}$} $\it 4.82(3)\times 10^{-4}$} $\it-7.637(3)\times 10^{-3}$} $\it-1.84018(8)\times 10^{-1}$}
$2p_{1/2}$ $52(2)\times 10^{-8} 1.1$$171(7)\times 10^{-7}$ $1.872(3)\times 10^{-6}$ $2.994(1)\times 10^{-5}$ $5156(2)\times 10^{-4} 1.1$$16946(7)\times 10^{-3}$ $1.87161(3)\times 10^{-2}$ $2.99777(1)\times 10^{-1}$
$\it 4.78(2)\times 10^{-8}$} $\it 5.76(7)\times 10^{-8}$} $\it-1.175(3)\times 10^{-6}$} $\it-2.716(1)\times 10^{-5}$} $\it-1.4530(2)\times 10^{-4}$} $\it-1.15208(7)\times 10^{-3}$} $\it-1.86466(3)\times 10^{-2}$} $\it-2.99499(1)\times 10^{-1}$}
$2p_{3/2}$ $2.42(7)\times 10^{-6}$ $6.35(6)\times 10^{-6}$ $2.79(1)\times 10^{-5}$ $1.350(2)\times 10^{-4}$ $3.894(4)\times 10^{-4}$ $1.8340(7)\times 10^{-3}$ $2.1389(2)\times 10^{-2}$ $3.10529(1)\times 10^{-1}$
$\it-2.38(3)\times 10^{-6}$} $\it-6.3(2)\times 10^{-6}$} $\it-2.79(3)\times 10^{-5}$} $\it-1.305(4)\times 10^{-4}$} $\it-3.810(7)\times 10^{-4}$} $\it-1.828(1)\times 10^{-3}$} $\it-2.1354(1)\times 10^{-2}$} $\it-3.10272(4)\times 10^{-1}$}
55 $1s_{1/2}$ $1.3\times 10^{-12}$ $6\times 10^{-13} 5.5$$5\times 10^{-11}$ $1.1(4)\times 10^{-9}$ $6.0(8)\times 10^{-9}$ $4.7(2)\times 10^{-8}$ $7.62(9)\times 10^{-7}$ $1.222(4)\times 10^{-5}$
$2s_{1/2}$ $4.610(2)\times 10^{-8}$ $3.2532(5)\times 10^{-7}$ $5.0616(2)\times 10^{-6}$ $8.54177(9)\times 10^{-5}$ $4.94906(2)\times 10^{-4}$ $4.249908(6)\times 10^{-3}$ $1.48846(1)\times 10^{-2}$ $1.65475(2)\times 10^{-1}$
$\it-9.149(2)\times 10^{-8}$} $\it-4.5153(5)\times 10^{-7}$} $\it-5.5688(2)\times 10^{-6}$} $\it-8.74875(9)\times 10^{-5}$} $\it-4.99773(2)\times 10^{-4}$} $\it-4.264815(6)\times 10^{-3}$} $\it-1.49319(1)\times 10^{-2}$} $\it 1.65378(4)\times 10^{-1}$}
$2p_{1/2}$ $6.305(9)\times 10^{-9}$ $1.738(2)\times 10^{-8}$ $6.719(10)\times 10^{-8}$ $2.320(4)\times 10^{-7}$ $3.838(8)\times 10^{-7}$ $1.65(2)\times 10^{-7}$ $2.3045(9)\times 10^{-5}$ $4.4800(4)\times 10^{-4}$
$\it 8.81(2)\times 10^{-9}$} $\it 2.434(2)\times 10^{-8}$} $\it 9.487(9)\times 10^{-8}$} $\it 3.433(4)\times 10^{-7}$} $\it 6.338(9)\times 10^{-7}$} $\it 5.30(2)\times 10^{-7}$} $\it-2.0267(9)\times 10^{-5}$} $\it-4.3689(4)\times 10^{-4}$}
$2p_{3/2}$ $1.9848(9)\times 10^{-8}$ $1.5361(2)\times 10^{-7}$ $2.49167(10)\times 10^{-6}$ $4.25045(4)\times 10^{-5}$ $2.468149(9)\times 10^{-4}$ $2.121595(3)\times 10^{-3}$ $7.39879(6)\times 10^{-3}$ $8.33945(10)\times 10^{-2}$
$\it 3.9834(9)\times 10^{-8}$} $\it 2.0919(2)\times 10^{-7}$} $\it 2.71512(10)\times 10^{-6}$} $\it 4.34194(4)\times 10^{-5}$} $\it 2.4906(3)\times 10^{-4}$} $\it 2.135(2)\times 10^{-3}$} $\it 7.60(7)\times 10^{-3}$} $\it-8.22(4)\times 10^{-2}$}
$10$ $1s_{1/2}$ $2.8\times 10^{-12}$ $7.5\times 10^{-12}$ $2.7\times 10^{-11}$ $4.9\times 10^{-11}$ $1.0\times 10^{-10}$ $2(1)\times 10^{-9}$ $4.4(6)\times 10^{-8}$ $7.5(2)\times 10^{-7}$
$2s_{1/2}$ $3.860(5)\times 10^{-8}$ $1.1441(9)\times 10^{-7}$ $5.236(1)\times 10^{-7}$ $3.0224(4)\times 10^{-6}$ $1.02470(10)\times 10^{-5}$ $5.9106(3)\times 10^{-5}$ $8.1443(1)\times 10^{-4}$ $1.275844(4)\times 10^{-2}$
$\it-4.998(5)\times 10^{-8}$} $\it-1.4591(9)\times 10^{-7}$} $\it-6.494(1)\times 10^{-7}$} $\it-3.5253(4)\times 10^{-6}$} $\it-1.13786(10)\times 10^{-5}$} $\it-6.2250(3)\times 10^{-5}$} $\it-8.2703(1)\times 10^{-4}$} $\it-1.280920(4)\times 10^{-2}$}
$2p_{1/2}$ $3.799(1)\times 10^{-8}$ $1.0965(3)\times 10^{-7}$ $4.4716(6)\times 10^{-7}$ $1.79706(2)\times 10^{-6}$ $4.03993(4)\times 10^{-6}$ $1.11540(1)\times 10^{-5}$ $4.32084(5)\times 10^{-5}$ $1.50089(2)\times 10^{-4}$
$\it 3.854(6)\times 10^{-8}$} $\it 1.116(4)\times 10^{-7}$} $\it 4.541(2)\times 10^{-7}$} $\it 1.8226(4)\times 10^{-6}$} $\it 4.1076(1)\times 10^{-6}$} $\it 1.1316(7)\times 10^{-5}$} $\it 4.3896(9)\times 10^{-5}$} $\it 1.5280(3)\times 10^{-4}$}
$2p_{3/2}$ $3.07(7)\times 10^{-10}$ $2.37(2)\times 10^{-9}$ $3.796(8)\times 10^{-8}$ $6.075(3)\times 10^{-7}$ $3.0764(7)\times 10^{-6}$ $2.3763(2)\times 10^{-5}$ $3.82188(8)\times 10^{-4}$ $6.24933(3)\times 10^{-3}$
$\it 5.292(7)\times 10^{-9}$} $\it 1.622(2)\times 10^{-8}$} $\it 9.334(8)\times 10^{-8}$} $\it 8.290(3)\times 10^{-7}$} $\it 3.5749(7)\times 10^{-6}$} $\it 2.5149(2)\times 10^{-5}$} $\it 3.87740(8)\times 10^{-4}$} $\it 6.27172(3)\times 10^{-3}$}
$30$ $1s_{1/2}$ $2.4\times 10^{-12}$ $6.6\times 10^{-12}$ $2.6\times 10^{-11}$ $1.0\times 10^{-10}$ $2.3\times 10^{-10}$ $6.2\times 10^{-10}$ $2.1\times 10^{-9}$ $1.6\times 10^{-9}$
$2s_{1/2}$ $1.06(1)\times 10^{-9}$ $8.2(5)\times 10^{-9}$ $1.397(1)\times 10^{-7}$ $2.8793(3)\times 10^{-6}$ $1.2856(1)\times 10^{-5}$ $2.6091(7)\times 10^{-5}$ $2.7323(3)\times 10^{-4}$ $2.52239(2)\times 10^{-3}$
$\it-6.6(1)\times 10^{-10}$} $\it 3.4(5)\times 10^{-9}$} $\it 1.205(1)\times 10^{-7}$} $\it 2.7992(3)\times 10^{-6}$} $\it 1.2673(1)\times 10^{-5}$} $\it 2.5623(7)\times 10^{-5}$} $\it-2.7479(2)\times 10^{-4}$} $\it-2.52805(3)\times 10^{-3}$}
$2p_{1/2}$ $1.051(3)\times 10^{-9}$ $8.2(2)\times 10^{-9}$ $1.398(7)\times 10^{-7}$ $2.8807(3)\times 10^{-6}$ $1.28638(3)\times 10^{-5}$ $2.6155(3)\times 10^{-5}$ $2.72201(10)\times 10^{-4}$ $2.50577(5)\times 10^{-3}$
$\it-4.87(4)\times 10^{-10}$} $\it-6.6(2)\times 10^{-9}$} $\it-1.333(7)\times 10^{-7}$} $\it-2.871(3)\times 10^{-6}$} $\it-1.234(1)\times 10^{-5}$} $\it-3.2(3)\times 10^{-5}$} $\it 2.6(1)\times 10^{-4}$} $\it 2.33(8)\times 10^{-3}$}
$2p_{3/2}$ $1.9\times 10^{-12}$ $7.2\times 10^{-12}$ $6(4)\times 10^{-11}$ $8(2)\times 10^{-10}$ $4.0(4)\times 10^{-9}$ $3.0(1)\times 10^{-8}$ $4.81(4)\times 10^{-7}$ $7.68(2)\times 10^{-6}$
$\it 5.45(4)\times 10^{-10}$} $\it 1.52(1)\times 10^{-9}$} $\it 6.10(4)\times 10^{-9}$} $\it 2.50(2)\times 10^{-8}$} $\it 5.83(4)\times 10^{-8}$} $\it 1.81(1)\times 10^{-7}$} $\it 1.085(4)\times 10^{-6}$} $\it 1.009(2)\times 10^{-5}$}
$60$ $1s_{1/2}$ $1.6\times 10^{-12}$ $4.4\times 10^{-12}$ $1.8\times 10^{-11}$ $7.0\times 10^{-11}$ $1.6\times 10^{-10}$ $4.4\times 10^{-10}$ $1.7\times 10^{-9}$ $6.6\times 10^{-9}$
$2s_{1/2}$ $8(8)\times 10^{-12}$ $4(2)\times 10^{-11}$ $5.0(9)\times 10^{-10}$ $7.5(3)\times 10^{-9}$ $3.73(8)\times 10^{-8}$ $2.85(2)\times 10^{-7}$ $4.557(8)\times 10^{-6}$ $7.324(3)\times 10^{-5}$
$\it-3.78(8)\times 10^{-10}$} $\it-1.03(2)\times 10^{-9}$} $\it-3.79(9)\times 10^{-9}$} $\it-9.7(3)\times 10^{-9}$} $\it-1.4(8)\times 10^{-9}$} $\it 1.78(2)\times 10^{-7}$} $\it 4.127(8)\times 10^{-6}$} $\it 7.152(3)\times 10^{-5}$}
$2p_{1/2}$ $2.4\times 10^{-12}$ $2.5(9)\times 10^{-11}$ $4.4(3)\times 10^{-10}$ $7.2(1)\times 10^{-9}$ $3.68(3)\times 10^{-8}$ $2.847(9)\times 10^{-7}$ $4.565(4)\times 10^{-6}$ $7.346(2)\times 10^{-5}$
$\it 1.22(3)\times 10^{-10}$} $\it 3.20(9)\times 10^{-10}$} $\it 9.4(3)\times 10^{-10}$} $\it-1.7(1)\times 10^{-9}$} $\it-2.44(3)\times 10^{-8}$} $\it-2.502(9)\times 10^{-7}$} $\it-4.428(4)\times 10^{-6}$} $\it-7.291(2)\times 10^{-5}$}
$2p_{3/2}$ $1.5\times 10^{-12}$ $4.2\times 10^{-12}$ $1.7\times 10^{-11}$ $7.5\times 10^{-11}$ $1.9\times 10^{-10}$ $7.5\times 10^{-10}$ $7(3)\times 10^{-9}$ $9(1)\times 10^{-8}$
$\it 1.28(3)\times 10^{-10}$} $\it 3.56(9)\times 10^{-10}$} $\it 1.42(3)\times 10^{-9}$} $\it 5.7(1)\times 10^{-9}$} $\it 1.28(3)\times 10^{-8}$} $\it 3.59(9)\times 10^{-8}$} $\it 1.48(3)\times 10^{-7}$} $\it 6.5(1)\times 10^{-7}$}
$92$ $1s_{1/2}$ $1.6\times 10^{-11}$ $4.5\times 10^{-11}$ $1.8\times 10^{-10}$ $7.1\times 10^{-10}$ $1.6\times 10^{-9}$ $4.5\times 10^{-9}$ $1.8\times 10^{-8}$ $7.1\times 10^{-8}$
$2s_{1/2}$ $5.4\times 10^{-14}$ $3.0\times 10^{-13}$ $4.0\times 10^{-12}$ $6.2\times 10^{-11}$ $3.1\times 10^{-10}$ $2.4\times 10^{-9}$ $4(3)\times 10^{-8}$ $6(1)\times 10^{-7}$
$\it-1.3(2)\times 10^{-10}$} $\it-3.6(7)\times 10^{-10}$} $\it-1.4(3)\times 10^{-9}$} $\it-6(1)\times 10^{-9}$} $\it-1.3(2)\times 10^{-8}$} $\it-3.4(7)\times 10^{-8}$} $\it-1.1(3)\times 10^{-7}$} $\it 2.4\times 10^{-8}$}
$2p_{1/2}$ $4.3\times 10^{-13}$ $1.0\times 10^{-12}$ $1.3\times 10^{-12}$ $4.1\times 10^{-11}$ $3(1)\times 10^{-10}$ $2.3(3)\times 10^{-9}$ $3.8(1)\times 10^{-8}$ $6.18(5)\times 10^{-7}$
$\it 4.0(1)\times 10^{-11}$} $\it 1.12(3)\times 10^{-10}$} $\it 4.4(1)\times 10^{-10}$} $\it 1.73(5)\times 10^{-9}$} $\it 3.7(1)\times 10^{-9}$} $\it 8.8(3)\times 10^{-9}$} $\it 6(1)\times 10^{-9}$} $\it-4.41(5)\times 10^{-7}$}
$2p_{3/2}$ $1.2\times 10^{-12}$ $3.5\times 10^{-12}$ $1.4\times 10^{-11}$ $5.6\times 10^{-11}$ $1.3\times 10^{-10}$ $3.5\times 10^{-10}$ $1.5\times 10^{-9}$ $7.7\times 10^{-9}$
$\it 4.7(3)\times 10^{-11}$} $\it 1.32(8)\times 10^{-10}$} $\it 5.3(3)\times 10^{-10}$} $\it 2.1(1)\times 10^{-9}$} $\it 4.7(3)\times 10^{-9}$} $\it 1.32(8)\times 10^{-8}$} $\it 5.3(3)\times 10^{-8}$} $\it 2.1(1)\times 10^{-7}$}
[…]
black Several observations can be drawn from Table 3. Considering the ground state of the hydrogen atom, a clear $T^{4}$ dependence is evident. This can be verified by comparing the values at temperatures of $10$ K and $100$ K, or $20$ K and $200$ K. This dependence holds approximately up to $Z=10$, beyond which a deviation from this asymptotic behavior is observed. For $Z=30$ and higher, the $T^{4}$ dependence is replaced by a quadratic one, which is explicitly demonstrated for $Z=92$. We attribute this to the fact that the $T^{4}$ estimate was derived within the non-relativistic limit, which, as the results in Table 3 demonstrate, is no longer sufficiently accurate for $Z>10$. Thus, relativistic thermal effects become fundamentally important for high nuclear charges.
[…]
black Another equally interesting effect is the deviation from the $Z^{-4}$ dependence with increasing nuclear charge at low temperatures. The breakdown of the non-relativistic $Z^{-4}$ scaling occurs earlier at temperatures below room temperature. For instance, this is clearly manifested for the ground state at temperatures of $50$ K and $10$ K for $Z=1$ and $Z=10$, respectively. We also note the sign-alternating behavior of the thermal shift with varying temperature and nuclear charge. Again, since this estimate was obtained in the non-relativistic limit, we attribute this behavior to the influence of relativistic effects. A discussion of the latter can be found in [26].
@@
Being the most accurately calculated to date for the hydrogen atom, the values for the thermal shift, presented in Table 2, can be used to determine the thermal frequency shifts in transitions between fine sublevels. In accordance with the quantum mechanical formalism for incorporating thermal radiation corrections to atomic levels [14], or via the QED one-loop perturbative framework [42, 26], the dominant thermal contribution arises due to the static dipole polarizability. Considering that in this case the perturbation operator does not act on the spin part of the electronic function, such contributions cancel out for states with equal principal quantum number $n$ and orbital angular momentum $l$, but different total angular momenta $j$.
@@
The detailed theoretical analysis assigning the influence of the "electric component"  (Stark effect) and the "magnetic component"  (Zeeman effect), as well as the derivation of other thermal corrections, has been presented quite recently in the work [26]. A distinctive feature of the approach presented in this study is that the fine structure of atomic levels is automatically taken into account within the framework of Dirac’s theory. Thus, by performing calculations with a sufficient number of significant digits, we can accurately account for the mentioned reduction.
@@
[…]
@@ H $\Re\Delta\varepsilon_{3s_{1/2}}^{\beta}-\Re\Delta\varepsilon_{1s_{1/2}}^{\beta}$ $T_{m}$
To perform the numerical calculations, see Eq. (13), the values for the hyperfine splitting were borrowed from Refs. [37, 19, 36, 30, 39]. First, what follows from Table 6 is the cubic decrease of the thermal frequency shift with an increase in the principal quantum number (easily verifiable for $s$-states). Secondly, it can be noted that the thermal shift actually increases cubically with the growth of the nuclear charge $Z$, as well as quadratically with the increase in temperature approaching laboratory conditions.
The corresponding estimate arises from formula (13), considering that at room temperature, the thermal radiation frequency $\omega$ is on the order of $2.82\beta\approx 0.00268$ in atomic units, while for hydrogen $\Delta\varepsilon_{\rm HFS}\approx 2.158\times 10^{-7}$ a.u. Then, expanding expression (13) in a series with respect to the small parameter $\Delta\varepsilon_{\rm HFS}$, one can obtain $\Delta\varepsilon_{\rm HFS}^{\beta}\sim\Delta\varepsilon_{\rm HFS}(k_{\rm B}T)^{2}$, $\Delta\varepsilon_{\rm HFS}\sim Z^{3}$ [37, 3]. However, this estimate is valid only up to $Z=23$. For larger values of the nuclear charge, it is necessary to perform a decomposition by the frequency of photons $\omega$ from thermal radiation. Then, a different dependence will emerge, and moreover, one of the opposite sign: $\Delta\varepsilon_{\rm HFS}^{\beta}\sim-(k_{\rm B}T)^{4}/\Delta\varepsilon_{\rm HFS}$. In the intermediate region, the complete form of equation (13) should be considered. In certain specific cases, the effect of finite lifetime can play a role [26].
@@
The value of the thermal shift to the hyperfine transition frequency in the hydrogen atom agrees well with the results of [20], but it is negligibly small even with respect to the accuracy of the experiment of [18]. However, for neutral and singly ionized atoms, the thermal shift of the hyperfine splitting remains the same in terms of the ratio of the shift to the transition frequency and undoubtedly makes a significant contribution to the budget of experimental error [45, 26]. Below we present a graph in Fig. 6 illustrating the $Z$-dependence of the shift $\Delta\varepsilon_{a}^{\beta}$ expressed by Eq. (13).
[…]
@@
Figure 7: The graph shows the dependence of the hyperfine splitting interval of the ground state in hydrogen-like ions on the nuclear charge $Z$ (in logarithmic scale). The points represent the used values, $\Delta\varepsilon_{\rm HFS}$ in eV. The dashed line corresponds to the dependence $\frac{1}{2}\Delta\varepsilon_{\rm HFS}^{Z=1}Z^{3}$, chosen as the interpolation according to [37], see Table 8 for details.
@@
[…]
@@
Combining the results from Tables 6, 7, and Fig. 6, one can conclude that the thermal shift for the hyperfine splitting frequency of the ground state (despite the cubic increase with $Z$) is unlikely to be related to corresponding measurements in highly charged ions, see, e.g., [3].